OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 14 and the last term is -3,919,104?

OpenStudy (anonymous):

the rule is ar^n-1=Xn

OpenStudy (anonymous):

for the term

OpenStudy (anonymous):

and a is the first term so it will be equal 14

OpenStudy (anonymous):

14r^n-1=14r^7=the last term

OpenStudy (anonymous):

then take the value of r and put it in the sum of series rule = a/1-r